# Telegrapher’s Generative Model via Kac Flows

### v4

@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
$u(t,x)=\frac{1}{2}\mathbb{E}\Big[f_{0}\big(x+S(t)\big)\Big]+\frac{1}{2}\mathbb{E}\Big[f_{0}\big(x-S(t)\big)\Big]$ (27)
@@
Further, if $f_{0}$ is a probability density of a random variable $X_{0}$ independent of $S(t)$, then (27) is the probability density of the Kac process $(X_{t})_{t\geq 0}$ starting in $X_{0}\sim f_{0}$ defined by (26).
@@
$\mathbb{E}\Big[f_{0}\big(x-S(t)\big)\Big]$ $=\int_{\Omega}f_{0}(x-S(t)(\omega)\,\mathrm{d}\mathbb{P}(\omega)=\int_{S(t)(\Omega)}f_{0}(x-y)\,\mathrm{d}P_{S(t)}(y)$ (28)
$=(f_{0}\ast P_{S(t)})(x)=p_{X_{0}+S(t)}(x),$ (29)
@@
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [29] and [25] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [34, Chapter 4.4].
Lemma 2.
@@
$\quad+\beta\int_{x-ct}^{x+ct}I_{0}(\beta r_{t}(x-y))\,f_{0}(y)\,\mathrm{d}y\Big),$ (31)
@@
$u(t,x)$ $=\frac{1}{2}e^{-at}\big(\delta_{0}(x+ct)+\delta_{0}(x-ct)\big)+\tilde{u}(t,x),$ (32)
@@
$\tilde{u}(t,x)$ $\coloneqq\frac{1}{2}e^{-at}\Big(\beta ct\frac{I_{0}^{\prime}(\beta r_{t}(x))}{r_{t}(x)}+\beta I_{0}(\beta r_{t}(x))\Big){1}_{[-ct,ct]}(x).$ (33)
@@
The measure $\mu_{t}$ is exactly the probability distribution of the Kac process $K(t)$ starting in 0 from (25), see [15, Section 2].
@@
Remark 3.
By [15, Section 2], the absolute continuous part of the distribution of the initially right-moving and initially left-moving Kac process $\pm S(t)$ starting in 0 is given by
$\tilde{u}_{\pm}(t,x)$ $=\frac{1}{2}e^{-at}\Big(\frac{\beta(ct\pm x)}{r_{t}(x)}\,I_{0}^{\prime}(\beta r_{t}(x))+\beta\,I_{0}(\beta r_{t}(x))\Big){1}_{[-ct,ct]}(x),$ (34)
@@
$\partial_{t}\tilde{u}_{+}+c\,\partial_{x}\tilde{u}_{+}=-a\,\tilde{u}_{+}+a\,\tilde{u}_{-},\quad\partial_{t}\tilde{u}_{-}-c\,\partial_{x}\tilde{u}_{-}=a\,\tilde{u}_{+}-a\,\tilde{u}_{-}$. (35)
Intuitively, the equations (35) can be understood as the continuity equation for the right-moving and left-moving particles including source and sink terms, see [32]. Adding the equations yields
@@
$\tilde{J}(t,x)$ $\coloneqq\frac{c}{2}\left(\tilde{u}_{+}(t,x)-\tilde{u}_{-}(t,x)\right)=\frac{1}{2}e^{-at}\,\frac{\beta cx}{r_{t}(x)}\,I_{0}^{\prime}(\beta r_{t}(x)),\quad x\in(-ct,ct).$ (36)
@@
$\partial_{t}\tilde{u}+\partial_{x}\tilde{J}=0.$ (37)
@@
After these preparations, we can explicitly describe the velocity field of the Kac process starting in a Dirac point which will be important for learning velocity fields later. The Kac velocity in (43) may be compared with those of coupling and diffusion flows given in (7) and (11), respectively.
Theorem 4.
Let $f_{0}=\delta_{0}$ be the Dirac measure on $\mathbb{R}$. Then $\mu_{t}$ associated to $u(t,\cdot)$ in (32) solves the continuity equation
$\partial_{t}\mu_{t}=-\partial_{x}(\mu_{t}v_{t})$ (38)
@@
$v(t,x)$ $\coloneqq\left\{\begin{array}[]{cl}\frac{x}{t+\frac{r_{t}(x)}{c}\frac{I_{0}(\beta r_{t}(x))}{I_{0}^{\prime}(\beta r_{t}(x))}}&\text{if}\quad x\in(-ct,ct),\\ \;\;\,c&\text{if}\quad x=ct,\\ -c&\text{if}\quad x=-ct,\\ \;\;\,\text{arbitrary}&\text{otherwise}.\end{array}\right$. (43)
@@
At this time, we will assume the existence of a velocity field $v_{t}$ satisfying the continuity equation (38) in a weak sense, and show that in this case $v_{t}$ must admit the form (43). Later, in Proposition 10, we generally and independently prove that $\mu_{t}$ is indeed absolutely continuous and admits a velocity field $v_{t}$.
Fix $t>0$. Formal differentiation of (32) leads to
$\partial_{t}u(t,x)=$ $-au(t,x)+\frac{1}{2}e^{-at}\big(c\delta_{0}^{\prime}(x+ct)-c\delta_{0}^{\prime}(x-ct)$ (44)
$+(a+a^{2}t)z_{1}(t,x)+atz_{2}(t,x)\big){1}_{[-ct,ct]}(x),$ (45)
@@
$z_{1}(t,x)\coloneqq\frac{I_{0}^{\prime}(\beta r_{t}(x))}{r_{t}(x)}\quad\text{and}\quad z_{2}(t,x)\coloneqq\frac{act\,I_{0}^{\prime\prime}(\beta r_{t}(x))-c^{2}t\,z_{1}(t,x)}{r_{t}(x)^{2}}.$ (46)
First, we consider the case $x\in(-ct,ct)$. Then, all the Dirac terms disappear and we only need to work with $\tilde{u}$ given in (33), and $\partial_{t}\tilde{u}$ is given in (44) without the Dirac terms. Using l’ Hospital’s rule and $I_{0}^{\prime}(0)=I_{0}^{(3)}(0)=0$, $I_{0}^{\prime\prime}(0)=\frac{1}{2}$, $I_{0}^{(4)}(0)=\frac{3}{8}$, it is easy to check that both functions $z_{1},z_{2}$ are bounded in $x\in(-ct,ct)$, so that $\partial_{t}\tilde{u}(t,\cdot)$ is integrable. Since $\tilde{u}(t,x)$ is symmetric with respect to $x$, the velocity $v(t,x)$ must be antisymmetric in $x$. By (38), we conclude
$2\tilde{u}(t,x)v(t,x)=-\int_{-x}^{x}\partial_{t}\tilde{u}(t,y)\,\mathrm{d}y\quad\text{ for all }x\in(-ct,ct).$ (47)
At the same time, we obtain from (37) and by the anti-symmetry of $\tilde{J}(t,\cdot)$ that
@@
Together with (47), (36) and (33), this implies
$v(t,x)=\frac{\tilde{J}(t,x)}{\tilde{u}(t,x)}=\frac{e^{-at}\,\frac{\beta cx}{r_{t}(x)}I_{0}^{\prime}(\beta r_{t}(x))}{e^{-at}\!\bigl[\frac{\beta ct}{r_{t}(x)}I_{0}^{\prime}(\beta r_{t}(x))+\beta I_{0}(\beta r_{t}(x))\bigr]}=\frac{x}{t+\frac{r_{t}(x)}{c}\,\frac{I_{0}(\beta r_{t}(x))}{I_{0}^{\prime}(\beta r_{t}(x))}},$ (49)
which shows (43) for all $x\in(-ct,ct)$.
@@
$\int_{-\infty}^{\infty}\partial_{t}u(t,x)\phi(x)\,\mathrm{d}x$ $=-\frac{ae^{-at}}{2}(\phi(-ct)+\phi(ct))+\frac{e^{-at}}{2}(-c\phi^{\prime}(-ct)+c\phi^{\prime}(ct))$ (50)
$\quad+\int_{-ct}^{ct}\partial_{t}\tilde{u}(t,x)\phi(x)\,\mathrm{d}x.$ (51)
On the other hand, the continuity equation (38) gives
$\int_{-\infty}^{\infty}\partial_{t}u(t,x)\phi(x)\,\mathrm{d}x$ $=\int_{-\infty}^{\infty}-\partial_{x}(u(t,x)v(t,x))\phi(x)\,\mathrm{d}x=\int_{-\infty}^{\infty}u(t,x)v(t,x)\phi^{\prime}(x)\,\mathrm{d}x$ (52)
$=\frac{e^{-at}}{2}(v(t,-ct)\phi^{\prime}(-ct)+v(t,ct)\phi^{\prime}(ct))+\int_{-ct}^{ct}\tilde{u}(t,x)v(t,x)\phi^{\prime}(x)\,\mathrm{d}x.$ (53)
@@
$\frac{e^{-at}}{2}(-c\cdot(-1)+c\cdot 1)=\frac{e^{-at}}{2}(-v(t,ct)\cdot(-1)+v(t,ct)\cdot 1),$ (54)
and hence, $v(t,ct)=c$. By anti-symmetry, it holds v(t,ct)=cv(t,-ct)=-c, and the proof of (43) is finished. ∎
@@ 5 Multi-dimensional Generalizations of the Telegrapher’s
The connection between the telegrapher’s equation and the Kac process can be generalized to higher dimensions via Kac’s insertion method which we briefly describe in Subsection 5.1. In contrast to one dimension, initial probability densities do not remain probability densities in the PDE formulation. Yet, we state a convergence result towards diffusion, which we will use subsequently. To circumvent the lack of conservation of mass, we propose a componentwise extension to higher dimensions in Subsection 5.2, and we prove regularity and convergence results of our componentwise process.
@@ 5.1 Kac’s Insertion Method
A generalization of Theorem 1 to multiple dimensions is given by the following theorem.
Theorem 5.
@@
$u(0,x)$ $=f_{0}(x),$ (58)
$\partial_{t}u(0,x)$ $=0.$ (59)
This claim was again conjectured in [17] and a proof was given in [11, Theorem 4] using semigroup theory.
Remark 6.
In 1D, the solution of (5) is given by $w_{c}(t,x)=\frac{1}{2}(f_{0}(x+ct)+f_{0}(x-ct))$, so that (57) reproduces (27). Unfortunately, in higher dimensions, the telegrapher’s solution (57) might not stay a probability density anymore. For example, the mass $\int_{\mathbb{R}^{3}}u(t,x)\,\mathrm{d}x$ in 3D behaves like $1-e^{-t}$, see [40, Ch. 2.2]. $\diamond$
Nevertheless, we will see that using an analogous insertion method for the diffusion, it is possible to show that the diffusion and telegrapher’s solutions converge to each other.
@@
Lemma 7.
For $f_{0}\in H^{2}(\mathbb{R}^{d})$, let $w$ be the solution of the undamped wave equation (5) with $c=1$. Let $W$ be a one-dimensional Wiener process. Then
$h(t,x):=\mathbb{E}\left[w\left(\sigma W(t),x\right)\right],\quad\sigma>0,$ (60)
@@
$\partial_{t}h(t,x)$ $=\frac{\sigma^{2}}{2}\Delta h(t,x),\quad x\in\mathbb{R}^{d},~t>0,$ (61)
$h(0,x)$ $=f_{0}(x).$ (62)
The statement can be found in [15], and a proof was given in [11, Theorem 5].
Note that if $f_{0}$ is a probability density and $\mathbf{X}_{0}\sim f_{0}$, then $h$ given by (60) is the probability density of a $d$-dimensional Wiener process $\sigma\mathbf{W}(t)$ starting in $\mathbf{X}_{0}$. Interestingly the functions $h$ and $u$ converge to each other for $t\to\infty$.
Theorem 8.
Fix $c_{0}>0$. For $f_{0}\in H^{2}(\mathbb{R}^{d})$, let $u$ be given by (57) with $a>0,\,c\geq c_{0}$, and $h$ by (60) with $\sigma^{2}:=\frac{c^{2}}{a}$. Then there exists a constant $C=C(c_{0},d)$ such that
$\|h(t,\cdot)-u(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq C\Big(\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\Big)\quad\text{ for all }t>0.$ (63)
@@
It is well-known that the Laplacian $\Delta:H^{2}(\mathbb{R}^{d})\subset L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ is a self-adjoint linear operator with non-positive spectrum $\sigma(\Delta)=(-\infty,0]$. By [23, Corollary 2] the solution of (5) with $c=1$ is given by $w(t,\cdot)=C(t)f_{0}$, where $(C(t))_{t\geq 0}$ is a uniformly bounded so-called cosine operator function with $\|C(t)\|_{L_{2}\to L_{2}}\leq 1$ for all $t\geq 0$. In particular, it follows that $\|w(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq\|f_{0}\|_{L_{2}(\mathbb{R}^{d})}$ is uniformly bounded in $t$. Now, [15, Theorem 2] yields the claim. ∎
Remark 9.
By (63), we see that for any $t\geq 0$, the solution $u^{a,c}(t,\cdot)$ of the telegrapher’s equation (5) converges to the solution $h(t,\cdot)$ of the diffusion equation (61) for $a,c\to\infty$ with fixed $\sigma^{2}=\frac{c^{2}}{a}$.  $\diamond$

### v5

@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
$u(t,x)=\frac{1}{2}\mathbb{E}\Big[f_{0}\big(x+S(t)\big)\Big]+\frac{1}{2}\mathbb{E}\Big[f_{0}\big(x-S(t)\big)\Big]$ (4.8)
@@
Further, if $f_{0}$ is a probability density of a random variable $X_{0}$ independent of $S(t)$, then (4.8) is the probability density of the Kac process $(X_{t})_{t\geq 0}$ starting in $X_{0}\sim f_{0}$ defined by (4.7).
@@
$\mathbb{E}\Big[f_{0}\big(x-S(t)\big)\Big]$ $=\int_{\Omega}f_{0}(x-S(t)(\omega)\,\mathrm{d}\mathbb{P}(\omega)=\int_{S(t)(\Omega)}f_{0}(x-y)\,\mathrm{d}P_{S(t)}(y)$ (4.9)
$=(f_{0}\ast P_{S(t)})(x)=p_{X_{0}+S(t)}(x),$ (4.10)
@@
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [34] and [31] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [38, Chapter 4.4].
Lemma 4.2.
@@
$\quad+\beta\int_{x-ct}^{x+ct}I_{0}(\beta r_{t}(x-y))\,f_{0}(y)\,\mathrm{d}y\Big),$ (4.12)
@@
$u(t,x)$ $=\frac{1}{2}e^{-at}\big(\delta_{0}(x+ct)+\delta_{0}(x-ct)\big)+\tilde{u}(t,x),$ (4.13)
@@
$\tilde{u}(t,x)$ $\coloneqq\frac{1}{2}e^{-at}\Big(\beta ct\frac{I_{0}^{\prime}(\beta r_{t}(x))}{r_{t}(x)}+\beta I_{0}(\beta r_{t}(x))\Big){1}_{[-ct,ct]}(x).$ (4.14)
@@
The measure $\mu_{t}$ is exactly the probability distribution of the Kac process $K(t)$ starting in 0 from (4.6), see [20, Section 2].
@@
Remark 4.3.
By [20, Section 2], the absolute continuous part of the distribution of the initially right-moving and initially left-moving Kac process $\pm S(t)$ starting in 0 is given by
$\tilde{u}_{\pm}(t,x)$ $=\frac{1}{2}e^{-at}\Big(\frac{\beta(ct\pm x)}{r_{t}(x)}\,I_{0}^{\prime}(\beta r_{t}(x))+\beta\,I_{0}(\beta r_{t}(x))\Big){1}_{[-ct,ct]}(x),$ (4.15)
@@
$\partial_{t}\tilde{u}_{+}+c\,\partial_{x}\tilde{u}_{+}=-a\,\tilde{u}_{+}+a\,\tilde{u}_{-},\quad\partial_{t}\tilde{u}_{-}-c\,\partial_{x}\tilde{u}_{-}=a\,\tilde{u}_{+}-a\,\tilde{u}_{-}$. (4.16)
Intuitively, the equations (4.16) can be understood as the continuity equation for the right-moving and left-moving particles including source and sink terms, see [36]. Adding the equations yields
@@
$\tilde{J}(t,x)$ $\coloneqq\frac{c}{2}\left(\tilde{u}_{+}(t,x)-\tilde{u}_{-}(t,x)\right)=\frac{1}{2}e^{-at}\,\frac{\beta cx}{r_{t}(x)}\,I_{0}^{\prime}(\beta r_{t}(x)),\quad x\in(-ct,ct).$ (4.17)
@@
$\partial_{t}\tilde{u}+\partial_{x}\tilde{J}=0.$ (4.18)
@@
After these preparations, we can explicitly describe the velocity field of the Kac process starting in a Dirac point which will be important for learning velocity fields later. The Kac velocity in (4.24) may be compared with those of coupling and diffusion flows given in (3.2) and (3.6), respectively.
Theorem 4.4.
Let $f_{0}=\delta_{0}$ be the Dirac measure on $\mathbb{R}$. Then $\mu_{t}$ associated to $u(t,\cdot)$ in (4.13) solves the continuity equation
$\partial_{t}\mu_{t}=-\partial_{x}(\mu_{t}v_{t})$ (4.19)
@@
$v(t,x)$ $\coloneqq\left\{\begin{array}[]{cl}\frac{x}{t+\frac{r_{t}(x)}{c}\frac{I_{0}(\beta r_{t}(x))}{I_{0}^{\prime}(\beta r_{t}(x))}}&\text{if}\quad x\in(-ct,ct),\\ \;\;\,c&\text{if}\quad x=ct,\\ -c&\text{if}\quad x=-ct,\\ \;\;\,\text{arbitrary}&\text{otherwise}.\end{array}\right$. (4.24)
@@
At this time, we will assume the existence of a velocity field $v_{t}$ satisfying the continuity equation (4.19) in a weak sense, and show that in this case $v_{t}$ must admit the form (4.24). Later, in Proposition 5.6, we generally and independently prove that $\mu_{t}$ is indeed absolutely continuous and admits a velocity field $v_{t}$.
Now, given some velocity field $v_{t}$ satisfying (4.19), we consider $\tilde{v}_{t}\coloneqq\frac{1}{2}(v_{t}(x)-v_{t}(-x))$, which is antisymmetric by definition. Using that $v_{t}$ solves (4.19) and the symmetry of $u(t,x)$ in $x$, it follows that also $\tilde{v}_{t}$ solves (4.19). Hence, in the following, we can assume the existence of some antisymmetric velocity field $v_{t}$ solving the continuity equation (4.19).
Fix $t>0$. Formal differentiation of (4.13) leads to
$\partial_{t}u(t,x)=$ $-au(t,x)+\frac{1}{2}e^{-at}\big(c\delta_{0}^{\prime}(x+ct)-c\delta_{0}^{\prime}(x-ct)$ (4.25)
$+(a+a^{2}t)z_{1}(t,x)+atz_{2}(t,x)\big){1}_{[-ct,ct]}(x),$ (4.26)
@@
$z_{1}(t,x)\coloneqq\frac{I_{0}^{\prime}(\beta r_{t}(x))}{r_{t}(x)}\quad\text{and}\quad z_{2}(t,x)\coloneqq\frac{act\,I_{0}^{\prime\prime}(\beta r_{t}(x))-c^{2}t\,z_{1}(t,x)}{r_{t}(x)^{2}}.$ (4.27)
First, we consider the case $x\in(-ct,ct)$. Then, all the Dirac terms disappear and we only need to work with $\tilde{u}$ given in (4.14), and $\partial_{t}\tilde{u}$ is given in (4.25) without the Dirac terms. Using l’ Hospital’s rule and $I_{0}^{\prime}(0)=I_{0}^{(3)}(0)=0$, $I_{0}^{\prime\prime}(0)=\frac{1}{2}$, $I_{0}^{(4)}(0)=\frac{3}{8}$, it is easy to check that both functions $z_{1},z_{2}$ are bounded in $x\in(-ct,ct)$, so that $\partial_{t}\tilde{u}(t,\cdot)$ is integrable. By our above arguments, the velocity field $v(t,x)$ can be assumed to be antisymmetric in $x$. Together with the symmetry of $\tilde{u}(t,x)$ in $x$, we conclude by (4.19) that
$2\tilde{u}(t,x)v(t,x)=-\int_{-x}^{x}\partial_{t}\tilde{u}(t,y)\,\mathrm{d}y\quad\text{ for all }x\in(-ct,ct).$ (4.28)
At the same time, we obtain from (4.18) and by the anti-symmetry of $\tilde{J}(t,\cdot)$ that
@@
Together with (4.28), (4.17) and (4.14), this implies
$v(t,x)=\frac{\tilde{J}(t,x)}{\tilde{u}(t,x)}=\frac{e^{-at}\,\frac{\beta cx}{r_{t}(x)}I_{0}^{\prime}(\beta r_{t}(x))}{e^{-at}\!\bigl[\frac{\beta ct}{r_{t}(x)}I_{0}^{\prime}(\beta r_{t}(x))+\beta I_{0}(\beta r_{t}(x))\bigr]}=\frac{x}{t+\frac{r_{t}(x)}{c}\,\frac{I_{0}(\beta r_{t}(x))}{I_{0}^{\prime}(\beta r_{t}(x))}},$ (4.30)
which shows (4.24) for all $x\in(-ct,ct)$.
@@
$\int_{-\infty}^{\infty}\partial_{t}u(t,x)\phi(x)\,\mathrm{d}x$ $=-\frac{ae^{-at}}{2}(\phi(-ct)+\phi(ct))+\frac{e^{-at}}{2}(-c\phi^{\prime}(-ct)+c\phi^{\prime}(ct))$ (4.31)
$\quad+\int_{-ct}^{ct}\partial_{t}\tilde{u}(t,x)\phi(x)\,\mathrm{d}x.$ (4.32)
On the other hand, the continuity equation (4.19) gives
$\int_{-\infty}^{\infty}\partial_{t}u(t,x)\phi(x)\,\mathrm{d}x$ $=\int_{-\infty}^{\infty}-\partial_{x}(u(t,x)v(t,x))\phi(x)\,\mathrm{d}x=\int_{-\infty}^{\infty}u(t,x)v(t,x)\phi^{\prime}(x)\,\mathrm{d}x$ (4.33)
$=\frac{e^{-at}}{2}(v(t,-ct)\phi^{\prime}(-ct)+v(t,ct)\phi^{\prime}(ct))+\int_{-ct}^{ct}\tilde{u}(t,x)v(t,x)\phi^{\prime}(x)\,\mathrm{d}x.$ (4.34)
@@
$\frac{e^{-at}}{2}(-c\cdot(-1)+c\cdot 1)=\frac{e^{-at}}{2}(-v(t,ct)\cdot(-1)+v(t,ct)\cdot 1),$ (4.35)
and hence, $v(t,ct)=c$. By anti-symmetry, it holds v(t,ct)=cv(t,-ct)=-c, and the proof of (4.24) is finished. ∎
@@ 5 Multi-dimensional Generalizations of the Telegrapher’s
The connection between the telegrapher’s equation and the Kac process can be generalized to higher dimensions via Kac’s insertion method which we briefly describe in Subsection 5.1. In contrast to one dimension, initial probability densities do not remain probability densities in the PDE formulation. Yet, we state a convergence result towards diffusion, which we will use subsequently. To circumvent the lack of conservation of mass, we propose two types of componentwise extensions to higher dimensions in Subsection 5.2, and we prove regularity and convergence results of our componentwise processes.
@@ 5.1 Kac’s Insertion Method
By inserting the random time (4.5) into the undamped wave solution, Theorem 4.1 can be generalized to multiple dimensions.
Theorem 5.1.
@@
$u(0,x)$ $=f_{0}(x),$ (5.4)
$\partial_{t}u(0,x)$ $=0.$ (5.5)
This claim was again conjectured in [22] and a proof was given in [14, Theorem 4] using semigroup theory. $\diamond$
The following observation states a defect of the multi-dimensional telegrapher’s equation (5.1) and crucially motivates our componentwise approach proposed in the next Subsection 5.2.
Remark 5.2.
In 1D, the solution of (5.1) is given by $w_{c}(t,x)=\frac{1}{2}(f_{0}(x+ct)+f_{0}(x-ct))$, so that (5.3) reproduces (4.8). Unfortunately, in higher dimensions, the telegrapher’s PDE solution (5.3) might not stay a probability density anymore. For example, the mass $\int_{\mathbb{R}^{3}}u(t,x)\,\mathrm{d}x$ in 3D behaves like $1-e^{-t}$, see [44, Ch. 2.2]. $\diamond$
Nevertheless, we will see that using an analogous insertion method for the diffusion, it is possible to show that the diffusion and telegrapher’s PDE solutions converge to each other.
@@
Lemma 5.3.
For $f_{0}\in H^{2}(\mathbb{R}^{d})$, let $w$ be the solution of the undamped wave equation (5.1) with $c=1$. Let $W$ be a one-dimensional Wiener process. Then
$h(t,x):=\mathbb{E}\left[w\left(\sigma W(t),x\right)\right],\quad\sigma>0,$ (5.6)
@@
$\partial_{t}h(t,x)$ $=\frac{\sigma^{2}}{2}\Delta h(t,x),\quad x\in\mathbb{R}^{d},~t>0,$ (5.7)
$h(0,x)$ $=f_{0}(x).$ (5.8)
The statement can be found in [20], and a proof was given in [14, Theorem 5]. $\diamond$
Note that if $f_{0}$ is a probability density and $\mathbf{X}_{0}\sim f_{0}$, then the solution $h$ given by (5.6) is the probability density of a $d$-dimensional Wiener process $\sigma\mathbf{W}(t)$ starting in $\mathbf{X}_{0}$.
Interestingly the functions $u$ and $h$ obtained by the insertion methods (5.3) and (5.6) converge to each other for $t\to\infty$. The proof of this fact crucially exploits the properties of the undamped wave solution (5.1).
Theorem 5.4.
Fix $c_{0}>0$. For $f_{0}\in H^{2}(\mathbb{R}^{d})$, let $u$ be given by (5.3) with $a>0,\,c\geq c_{0}$, and $h$ by (5.6) with $\sigma^{2}:=\frac{c^{2}}{a}$. Then there exists a constant $C=C(c_{0},d)$ such that
$\|h(t,\cdot)-u(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq C\Big(\frac{1}{at}+\frac{1}{c(ta)^{\frac{1}{2}}}\Big)\quad\text{ for all }t>0.$ (5.9)
@@
It is well-known that the Laplacian $\Delta:H^{2}(\mathbb{R}^{d})\subset L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ is a self-adjoint linear operator with non-positive spectrum $\sigma(\Delta)=(-\infty,0]$. By [29, Corollary 2] the solution of (5.1) with $c=1$ is given by $w(t,\cdot)=C(t)f_{0}$, where $(C(t))_{t\geq 0}$ is a uniformly bounded so-called cosine operator function with $\|C(t)\|_{L_{2}\to L_{2}}\leq 1$ for all $t\geq 0$. In particular, it follows that $\|w(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq\|f_{0}\|_{L_{2}(\mathbb{R}^{d})}$ is uniformly bounded in $t$. Now, [20, Theorem 2] yields the claim. ∎
Remark $5.$
Combining Theorems 5.1, 5.4 and Lemma 5.3, we see that for any $t\geq 0$, the solution $u^{a,c}(t,\cdot)$ of the telegrapher’s equation (5.1) converges to the solution $h(t,\cdot)$ of the diffusion equation (5.7) for $a,c\to\infty$ with fixed $\sigma^{2}=\frac{c^{2}}{a}$.  $\diamond$